OpenStudy (cookiemonster2412):

Solve the following right triangles. (Find all of the missing sides and angles.)

OpenStudy (cookiemonster2412):



OpenStudy (cookiemonster2412):

it's 50 degrees

jhonyy9 (jhonyy9):

so you know that sin 50 = 21 /VD VD = 21/sin 50 = 21/(0,26) = 80,7 VD=80,7 m<V = 90-50=40 m<V = 40 sin 40 = MD/80,7 MD = 80,7 *sin 40 MD = 80,7 *0,74 = 59,7 MD = 59,7 hope this will help you good luck bye

OpenStudy (cookiemonster2412):

thank you. it did help :)

jhonyy9 (jhonyy9):

was my pleasure bye